# The Orbits of Isolated Dwarfs in the Local Group from New 3D Kinematics: Constraints on First Infall, Backsplash, and Quenching Mechanisms

### v1

@@ 4.3 Leo T
Leo T is the lowest stellar mass galaxy with a confirmed HI reservoir. This HI reservoir is symmetric, with a low velocity dispersion disk (Adams & Oosterloo, 2018), supporting a history of non-interaction. Its SFH shows relatively constant star formation over the past $\sim$8 Gyrs (Weisz et al., 2012; Clementini et al., 2012). These studies also show possible cessation of star formation in Leo T in the past $\sim$25 Myrs, though this may be caused by the stochastic sampling of the initial mass function in such a low-mass galaxy (see Figure 5).
@@ 4.4 Pegasus
Pegasus’ recent ($\leq$6 Gyr) SFH shows an approximately constant star formation rate (Savino et al., 2025). This constant star formation rate also matches those of isolated dwarfs of Pegasus’ mass in simulations (Garrison-Kimmel et al., 2019). Our orbits substantiate these ideas, showing a very slim (0.8%) chance of Pegasus being a backsplash galaxy despite previous kinematic studies suggesting this was possible (Buck et al., 2019).
HI imaging of Pegasus shows an undisturbed solid-body-like HI disk (Kniazev et al., 2009) whose rotation is consistent with its RGB stars (Higgs et al., 2021), as expected for a classical dwarf that has had no interaction with a massive host.
@@ 4.5 Pisces
Pisces is the only dwarf in the present sample that resides within the virial radius of the MW or M31. At present, it is just inside of M31’s virial radius, having arrived there in just the last $\sim$100 Myr. Pisces has a small, disordered HI reservoir (Hunter et al., 2012) and an SFH that shows significant mass assembly in the early Universe, followed by a slow decline for most of the past 10 Gyrs (Hidalgo et al., 2011). The exception to this is a large burst of star formation about 1.5 Gyrs ago; this burst does not correlate with any encounter with M31 based on our orbital histories.
We concluded Pisces (LGS 3) is likely to be on first infall into M31’s halo, however, there has been much discussion historically about whether Pisces is a member of the M31 or M33 systems (Lee, 1995; Aparicio et al., 1997; Cook et al., 1999; McConnachie et al., 2005; McConnachie, 2012).
Recent RR Lyrae distance estimates (Savino et al., 2022) move Pisces significantly closer ($\sim$150 kpc) to the MW compared to previous TRGB estimates (McConnachie, 2012). The new 3D separation between Pisces and M31 (260 kpc) is similar to that between Pisces and M33 (235 kpc). In a companion paper, we will explore the recent interaction history between Pisces and M33 in the last $\sim$2 Gyr and how the categorization of Pisces as an associated dwarf of M31 versus M33 impacts interpretations of recent observational surveys.
@@ 4.6 Tucana
Tucana is commonly assumed to be a backsplash galaxy based on its SFH and HI gas properties. It has no detectable HI reservoir (Spekkens et al., 2014; Putman et al., 2021), and has no star formation in the past 8 Gyrs after it quenched around this time (Monelli et al., 2010b). Fu et al. (2024) also examined the metallicity distribution function of Tucana and compared it to other dwarfs in the LG. This showed that the metallicity gradient in Tucana is far more similar to those of the isolated galaxies of the LG (e.g., Leo A) rather than those that have interacted with the MW or M31 (e.g., Draco dSph).
@@
Simulated dwarfs with masses similar to that of Tucana ($1.1\times 10^{6}\,\rm M_{\odot}$) suggest these galaxies should be quenched even if isolated from the influence of a massive host (Digby et al., 2019; Garrison-Kimmel et al., 2019; Joshi et al., 2021). However, Tucana’s relatively low mass, compared to other dwarfs in the LG, lies in the regime where many zoom-in simulations suffer from resolution effects. The consequence of this is typically early quenching in almost all small dwarfs ($M_{*}\lesssim 1\times 10^{6}\,\rm M_{\odot}$; Garrison-Kimmel et al. 2019). This is in contrast to the LG, where dwarfs, such as Phoenix and Leo T, are still star-forming, despite having masses less than or equal to Tucana (Hidalgo et al., 2009; Weisz et al., 2012). Thus, most simulated dwarfs and Tucana quench for very different reasons.
Recently, the Lyra simulations (Gutcke et al., 2022), which simulated single isolated dwarfs to a mass resolution of individual stars, show that Tucana-mass dwarfs are at the boundary of being quenched by reionization. Some dwarfs in this mass range survive and continue to form stars until recent times, while others quench during reionization. However, these detailed simulations do not include any direct analogs for Tucana, where the dwarf survives reionization and then quenches later at intermediate times ($\sim$8 Gyrs).
Therefore, the evolution of Tucana is hard to explain with our current models of dwarf galaxy evolution. Studies of other isolated, quenched dwarfs such as Hedgehog (Li et al., 2024) also speculate about whether Hedgehog is a backsplash galaxy of the Centaurus A group. However, they also suggest other possible explanations for its quenching, including ram pressure stripping in the cosmic web (see also Benavides et al., 2025), reionization, or internal processes such as supernova and stellar feedback. The same mechanisms may play a major role in the quenching of Tucana. Observations of additional quenched field dwarfs are needed to disentangle these possible scenarios.
@@ 5 Summary and Conclusion
Star formation properties of the satellites in the SAGA survey, which took a census of satellite luminosity functions down to the magnitude of Leo I around 101 MW analogs, also defined as those within the projected virial radii of their host galaxies (Mao et al., 2024; Geha et al., 2024). (Mao et al., 2024) shows that the majority are star-forming, particularly far from the host galaxies. However, we do see that $\sim$20% are quenched even when at the virial radius of the host galaxy. This is also found in the ELVES sample (Carlsten et al., 2022). This implies that these dwarfs have quenched prior to encounters with the host or are backsplash galaxies.
If we examine the LG as a whole, we find 10 dwarfs isolated from either the MW or M31 in this mass range. Among these, Cetus and Tucana are the only dwarfs that are quenched, with the other 8 being star-forming dwarf irregulars or transition dwarfs. This is consistent with the quenched dwarfs in SAGA and ELVES at the edge of the host’s virial radius as being analogs to Cetus and Tucana, and therefore potentially quench prior to interaction with their hosts. Similarly, simulations (e.g., Garrison-Kimmel et al., 2019; Joshi et al., 2021; Samuel et al., 2022) typically quench dwarfs like Cetus and Tucana via host interactions, but this study’s results suggest alternative or earlier quenching mechanisms must be considered (e.g., reionization, internal feedback, dwarf-dwarf mergers).
@@
For example, the halos of all galaxies included in the orbital models are rigid, and therefore any deformation of the dark matter halos of the MW and M31, in particular, due to the passage of satellites is not included (e.g., Garavito-Camargo et al., 2021). As shown in Garavito-Camargo et al. (2021), the outskirts of the MW’s halo become deformed out to the virial radius. In forthcoming work (Patel, Garavito-Camargo et al., in prep.), we will explore the impact of this deformation on the orbits of LG dwarf galaxies.

### v2

@@ 4.3 Leo T
Leo T is the lowest stellar mass galaxy with a confirmed HI reservoir. This HI reservoir is symmetric, with a low velocity dispersion disk (E. A. K. Adams & T. A. Oosterloo, 2018), supporting a history of non-interaction. Its SFH shows relatively constant star formation over the past $\sim$8 Gyrs (D. R. Weisz et al., 2012; G. Clementini et al., 2012). These studies also show possible cessation of star formation in Leo T in the past $\sim$25 Myrs, though this may be caused by the stochastic sampling of the initial mass function in such a low-mass galaxy (see Figure 5).
@@ 4.4 Pegasus
Pegasus’ recent ($\leq$6 Gyr) SFH shows an approximately constant star formation rate (A. Savino et al., 2025). This constant star formation rate also matches those of isolated dwarfs of Pegasus’ mass in simulations (S. Garrison-Kimmel et al., 2019). Our orbits substantiate these ideas, showing a very slim (0.8%) chance of Pegasus being a backsplash galaxy despite previous kinematic studies suggesting this was possible (T. Buck et al., 2019).
HI imaging of Pegasus shows an undisturbed solid-body-like HI disk (A. Y. Kniazev et al., 2009) whose rotation is consistent with its RGB stars (C. R. Higgs et al., 2021), as expected for a classical dwarf that has had no interaction with a massive host.
@@ 4.5 Pisces
Pisces is the only dwarf in the present sample that resides within the virial radius of the MW or M31. At present, it is just inside of M31’s virial radius, having arrived there in just the last $\sim$100 Myr. Pisces has a small, disordered HI reservoir (D. A. Hunter et al., 2012) and an SFH that shows significant mass assembly in the early Universe, followed by a slow decline for most of the past 10 Gyrs (S. L. Hidalgo et al., 2011). The exception to this is a large burst of star formation about 1.5 Gyrs ago; this burst does not correlate with any encounter with M31 based on our orbital histories.
We concluded Pisces (LGS 3) is likely to be on first infall into M31’s halo, however, there has been much discussion historically about whether Pisces is a member of the M31 or M33 systems (M. G. Lee, 1995; A. Aparicio et al., 1997; K. H. Cook et al., 1999; A. W. McConnachie et al., 2005; A. W. McConnachie, 2012).
Recent RR Lyrae distance estimates (A. Savino et al., 2022) move Pisces significantly closer ($\sim$150 kpc) to the MW compared to previous TRGB estimates (A. W. McConnachie, 2012). The new 3D separation between Pisces and M31 (260 kpc) is similar to that between Pisces and M33 (235 kpc). In a companion paper, we will explore the recent interaction history between Pisces and M33 in the last $\sim$2 Gyr and how the categorization of Pisces as an associated dwarf of M31 versus M33 impacts interpretations of recent observational surveys.
@@ 4.6 Tucana
Tucana is commonly assumed to be a backsplash galaxy based on its SFH and HI gas properties. It has no detectable HI reservoir (K. Spekkens et al., 2014; M. E. Putman et al., 2021), and has no star formation in the past 8 Gyrs after it quenched around this time (M. Monelli et al., 2010b). S. W. Fu et al. (2024) also examined the metallicity distribution function of Tucana and compared it to other dwarfs in the LG. This showed that the metallicity gradient in Tucana is far more similar to those of the isolated galaxies of the LG (e.g., Leo A) rather than those that have interacted with the MW or M31 (e.g., Draco dSph).
@@
Simulated dwarfs with masses similar to that of Tucana ($1.1\times 10^{6}\,\rm M_{\odot}$) suggest these galaxies should be quenched even if isolated from the influence of a massive host (R. Digby et al., 2019; S. Garrison-Kimmel et al., 2019; G. D. Joshi et al., 2021). However, Tucana’s relatively low mass, compared to other dwarfs in the LG, lies in the regime where many zoom-in simulations suffer from resolution effects. The consequence of this is typically early quenching in almost all small dwarfs ($M_{*}\lesssim 1\times 10^{6}\,\rm M_{\odot}$; S. Garrison-Kimmel et al. 2019). This is in contrast to the LG, where dwarfs, such as Phoenix and Leo T, are still star-forming, despite having masses less than or equal to Tucana (S. L. Hidalgo et al., 2009; D. R. Weisz et al., 2012). Thus, most simulated dwarfs and Tucana quench for very different reasons.
Recently, the Lyra simulations (T. A. Gutcke et al., 2022), which simulated single isolated dwarfs to a mass resolution of individual stars, show that Tucana-mass dwarfs are at the boundary of being quenched by reionization. Some dwarfs in this mass range survive and continue to form stars until recent times, while others quench during reionization. However, these detailed simulations do not include any direct analogs for Tucana, where the dwarf survives reionization and then quenches later at intermediate times ($\sim$8 Gyrs).
Therefore, the evolution of Tucana is hard to explain with our current models of dwarf galaxy evolution. Studies of other isolated, quenched dwarfs such as Hedgehog (J. Li et al., 2024) also speculate about whether Hedgehog is a backsplash galaxy of the Centaurus A group. However, they also suggest other possible explanations for its quenching, including ram pressure stripping in the cosmic web (see also J. A. Benavides et al., 2025), reionization, or internal processes such as supernova and stellar feedback. The same mechanisms may play a major role in the quenching of Tucana. Observations of additional quenched field dwarfs are needed to disentangle these possible scenarios.
@@ 5 Summary and Conclusion
Star formation properties of the satellites in the SAGA survey, which took a census of satellite luminosity functions down to the magnitude of Leo I around 101 MW analogs, also defined as those within the projected virial radii of their host galaxies (Y.-Y. Mao et al., 2024; M. Geha et al., 2024). (Y.-Y. Mao et al., 2024) shows that the majority are star-forming, particularly far from the host galaxies. However, we do see that $\sim$20% are quenched even when at the virial radius of the host galaxy. This is also found in the ELVES sample (S. G. Carlsten et al., 2022). This implies that these dwarfs have quenched prior to encounters with the host or are backsplash galaxies.
If we examine the LG as a whole, we find 10 dwarfs isolated from either the MW or M31 in this mass range. Among these, Cetus and Tucana are the only dwarfs that are quenched, with the other 8 being star-forming dwarf irregulars or transition dwarfs. This is consistent with the quenched dwarfs in SAGA and ELVES at the edge of the host’s virial radius as being analogs to Cetus and Tucana, and therefore potentially quench prior to interaction with their hosts. Similarly, simulations (e.g., S. Garrison-Kimmel et al., 2019; G. D. Joshi et al., 2021; J. Samuel et al., 2022) typically quench dwarfs like Cetus and Tucana via host interactions, but this study’s results suggest alternative or earlier quenching mechanisms must be considered (e.g., reionization, internal feedback, dwarf-dwarf mergers).
@@
For example, the halos of all galaxies included in the orbital models are rigid, and therefore any deformation of the dark matter halos of the MW and M31, in particular, due to the passage of satellites is not included (e.g., N. Garavito-Camargo et al., 2021). As shown in N. Garavito-Camargo et al. (2021), the outskirts of the MW’s halo become deformed out to the virial radius. In forthcoming work (Patel, Garavito-Camargo et al., in prep.), we will explore the impact of this deformation on the orbits of LG dwarf galaxies.